Given that y varies directly with x, use the specified values to write a direct variation equation that relates x and y.
x=24, y=3

I don't get the instructions.
Thank you!
-Jenna

Answer by jim_thompson5910(35256)

"y varies directly with x" means that y = kx for some constant number k


Because x=24, y=3 , we can say that


y = kx

3 = k(24)

3/24 = k

1/8 = k

k = 1/8


So go back to the original equation y = kx and plug in k = 1/8 to get




So the answer is


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim